Would you please try to explain how to solve this problem: Determine the difference of the Surface Area of this objects to the nearest tenth of square foot: Object1: Rectangular Prism 10 by 5 by 5 Ft and 0bject 2: A cylinder 5 by 10 Ft.
Answer by ankor@dixie-net.com(22740)

Determine the difference of the Surface Area of this objects to the nearest tenth of square foot: Object1: Rectangular Prism 10 by 5 by 5 Ft and 0bject 2: A cylinder 5 by 10 Ft.
:
Rectangular prism surface area - cylinder surface area (including ends)
[2(L*W)+2(L*H)+2(W*H)] - [(pi*d*h) + (2pi*r^2)]
:
Diameter of cylinder given as 5 ft, therefore; Radius = 5/2 = 2.5 ft
diff = [2(10*5)+2(10*5)+2(5*5)] - [(pi*5*10) + (2pi*2.5^2)]
:
d = 2(50) + 2(50) + 2(25) - [157.08 + 39.27]
:
d = 250 - 196.35
:
d = 53.65, the difference in the surface areas